# Efficient Multimodal Dataset Distillation via Generative Models

### v1

@@ 1 Introduction
Dataset distillation wang2018datasetdistillation seeks to synthesize a compact dataset from a larger one, enabling a model trained on the distilled dataset to achieve strong performance on the original dataset. Typical kernel-based dataset distillation wang2018datasetdistillation ; zhou2022datasetfrepo aims to match the performance of the synthetic dataset with the original dataset. The matching training trajectories (MTT) approaches cazenavette2022datasetparameter_multi_1 ; li2023datasetmulti_2 ; cui2023scaling_multi_3 ; du2023minimizing_multi_4 ; zhao2024distillingltdd ; guo2023towardsdatm have been proven effective for the distillation of small datasets. Some works guo2023towardsdatm ; zhao2024distillingltdd even achieve a lossless performance from the original dataset. Recently, aiming the large-scale dataset, decoupled dataset distillation methods sun2024diversity ; yin2024squeeze , and generative distillation methods su2024d4m ; gu2024efficientminimax ; wang2025cao have been proposed, which enhances scalability by improving the distillation efficiency.
@@ 2.1 Dataset Distillation
Dataset distillation (DD) wang2018datasetdistillation aims to compress a large dataset into a small synthetic dataset while preserving the performance of the models trained on it. Dataset distillation methods can be categorized into several classes: performance matching wang2018datasetdistillation ; zhou2022datasetfrepo , parameter matching zhao2020datasetparameter_1 ; lee2022datasetparameter_2 ; jiang2023delvingparameter_3 ; cazenavette2022datasetparameter_multi_1 ; cui2023scaling_multi_3 ; zhao2024distillingltdd ; guo2023towardsdatm ; li2024prioritize ; wang2025emphasizing , distribution matching zhao2023dataset_distribution_1 ; wang2022cafe_distribution_2 ; shang2024mim4dd and decoupled distillation methods sun2024diversity ; yin2024squeeze . Recently, generative distillation methods gu2024efficientminimax ; su2024d4m ; wang2025cao synthesize images using generative models to efficiently distill large-scale datasets. Minimax Diffusion gu2024efficientminimax introduces additional minimax criteria in the generative training to enhance the performance of the generated images of the diffusion model. D4su2024d4m has explored the use of diffusion models to improve cross-architecture generalization and reduce computational overhead by constraining the consistency of the real and synthetic image spaces.

### v2

@@ 1 Introduction
Dataset distillation wang2018datasetdistillation seeks to synthesize a compact dataset from a larger one, enabling a model trained on the distilled dataset to achieve strong performance on the original dataset. Typical kernel-based dataset distillation wang2018datasetdistillation ; zhou2022datasetfrepo aims to match the performance of the synthetic dataset with the original dataset. The matching training trajectories (MTT) approaches cazenavette2022datasetmtt ; li2023datasetmulti_2 ; cui2023scalingtesla ; du2023minimizing_multi_4 ; zhao2024distillingltdd ; guo2023towardsdatm have been proven effective for the distillation of small datasets. Some works guo2023towardsdatm ; zhao2024distillingltdd even achieve a lossless performance from the original dataset. Recently, aiming the large-scale dataset, decoupled dataset distillation methods sun2024diversity ; yin2024squeeze , and generative distillation methods su2024d4m ; gu2024efficientminimax ; wang2025cao have been proposed, which enhances scalability by improving the distillation efficiency.
@@ 2.1 Dataset Distillation
Dataset distillation (DD) wang2018datasetdistillation aims to compress a large dataset into a small synthetic dataset while preserving the performance of the models trained on it. Dataset distillation methods can be categorized into several classes: performance matching wang2018datasetdistillation ; zhou2022datasetfrepo , parameter matching zhao2020datasetparameter_1 ; lee2022datasetparameter_2 ; jiang2023delvingparameter_3 ; cazenavette2022datasetmtt ; cui2023scalingtesla ; zhao2024distillingltdd ; guo2023towardsdatm ; li2024prioritize ; wang2025emphasizing , distribution matching zhao2023dataset_distribution_1 ; wang2022cafe_distribution_2 ; shang2024mim4dd and decoupled distillation methods sun2024diversity ; yin2024squeeze . Recently, generative distillation methods gu2024efficientminimax ; su2024d4m ; wang2025cao synthesize images using generative models to efficiently distill large-scale datasets. Minimax Diffusion gu2024efficientminimax introduces additional minimax criteria in the generative training to enhance the performance of the generated images of the diffusion model. D4su2024d4m has explored the use of diffusion models to improve cross-architecture generalization and reduce computational overhead by constraining the consistency of the real and synthetic image spaces.